Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $119,600 of manufacturing overhead for an estimated allocation base of $92,000 direct material dollars to be used in production.

The company has provided the following data for the just-completed year: Purchase of raw materials $ 132,000 Direct labor cost $ 88,000 Manufacturing overhead costs: Indirect labor $ 112,100 Property taxes $ 8,200 Depreciation of equipment $ 18,000 Maintenance $ 16,000 Insurance $ 9,700 Rent, building $ 31,000 Beginning Ending Raw Materials $ 26,000(Begining) $ 11,000 (Ending) Work in Process $ 47,000(Begining) $ 37,000 (Ending) Finished Goods $ 72,000 (Begining) $ 57,000 (Ending)

1. Compute the predetermined overhead rate for the year. 130%

2. Compute the amount of underapplied or overapplied overhead for the year. I got 191,100 (But it says it is wrong)

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $37,000 ending balance in Work in Process includes $8,200 of direct materials. Given this assumption, supply the information missing below:
Answer #1

1.

Predetermined overhead rate = Estimated overhead / Estimated direct material used

Predetermined overhead rate = $119,600 / 92,000 = 130%

2.

Over or underapplied overhead = Applied overhead - Actual overhead

Actual overhead = $112,100+8,200+18,000+16,000+9,700+31,000 = $195,000

Direct material used = $26,000+132,000-11,000 = $147,000

Applied overhead = $147,000 * 130% = $191,100

Over or underapplied overhead = $191,100 - 195,000

Underapplied overhead = $3,900

3.

Statement of Cost of Goods Manufactured
Beginning raw material $26,000
Raw material purchases 132,000
Ending raw material (11,000)
Direct material used in production 147,000
Direct labor 88,000
Manufacturing overhead applied 191,100
Total manufacturing cost 426,100
Beginning work in process 47,000
Ending work in process (37,000)
Cost of goods manufactured $436,100

4.

Unadjusted cost of goods sold = Beginning finished goods + Cost of goods manufactured - Ending finished goods

Unadjusted cost of goods sold = $72,000+436,100-57,000 = $451,100

5.

Direct material $8,200
Direct labor ($37,000-8,200-10,660) 18,140
Manufacturing overhead applied ($8,200*130%) 10,660
Ending work in process $37,000